OpenStudy (anonymous):

If sin theta = and cos theta < 0, what is tan theta ?

OpenStudy (misssunshinexxoxo):

What does sin theta equal? Might of got cut out?

OpenStudy (anonymous):

12/13

geerky42 (geerky42):

Ok sine is positive and cosine is negative. What quadrant is that?

OpenStudy (misssunshinexxoxo):

5-12-13 right triangle in the 4th quadrant. Draw it; label it and read the function values from the triangle. Short leg 5 is on the horizontal axis. Sine = LEG/HYP So 13² – 12² = 169 – 144 = 25 = 5² so the third leg is 5. The negative sign tells you it is in the 4th quadrant tan θ = –12/5 and cos θ = 5/13 and cot θ –5/12

OpenStudy (anonymous):

i know the answer i had a blonde moment i did the problem before and just went blank